# Mirror Descent Policy Optimisation for Robust Constrained Markov Decision Processes

### v2

@@ Appendix E Algorithm implementation details
All the algorithms were implemented in Pytorch. The code for RMCPMD was taken from its original implementation https://github.com/JerrisonWang/JMLR-DRPMD and the code for PPO was taken from the StableBaselines3 class. Both were modified to fit our purposes by allowing to turn off and on robust training and constraint-satisfaction. Similar to PPO, MDPO was also implemented following StableBaselines3 class conventions. All the environments are implemented in Gymnasium. For simplicity, the updates with LTMA are based on the TMA code in https://github.com/JerrisonWang/JMLR-DRPMD, which performs pure Monte Carlo based TMA rather than additional function approximation as proposed in our Approximate TMA algorithm. For MDPO and PPO experiments, we use four parallel environments and with four CPUs. Our longest experiments typically take no more than two hours to complete. The remainder of the section describes domain-specific hyperparameter settings.
@@ E.1 Hyperparameters for Cartpole experiments
Hyperparameters settings for the Carpole experiments can be found in 8. Settings for robust optimisation, including the policy architecture, transition kernel architecture, and TMA learning rate are taken from Wang et al., (2023) with the exception that we formulate a more challenging uncertainty set with a 5 times larger range. The policy learning rate and GAE lambda is typical for PPO based methods so we apply these for PPO and MDPO methods. The number of policy epochs and early stopping KL target is based for PPO methods on the standard repository for PPO-Lag (https://github.com/openai/safety-starter-agents/) and for MDPO it is based on the original paper’s settings (Tomar et al.,, 2022). PPO obtained better results without entropy regularisation and value function clipping on initial experiments, so for simplicity we disabled these for all algorithms. The batch size, multiplier initialisation, and multiplier learning rate and scaling are obtained via a limited tuning procedure. For the tuning, the batch size tried was in $\{100,200,400,1000,2000\}$ and the multiplier learning rate was set to $\{1\mathrm{e}^{-1},1\mathrm{e}^{-2},1\mathrm{e}^{-3},1\mathrm{e}^{-4}\}$ on partial runs. Then the use of a softplus transformation of the multiplier was compared to the direct (linear) multiplier, and for the linear multiplier we tested initialisations in $\{1,5,10\}$.
@@ E.2 Hyperparameters for Inventory Management experiments
Hyperparameters settings for the unidimensional Inventory Management (see Table 9) are similar to the Cartpole experiments. The key differences are the output layer, the discount factor, the transition kernel, and the TMA parameters, which follow settings of the domain in Wang et al., (2024). Based on initial tuning experiments with PPO and MDPO, the learning rates are set to $\eta=1\mathrm{e}^{-3}$ and $\eta_{\lambda}=1\mathrm{e}^{-2}$, a lower GAE lambda of 0.50 was chosen and PPO is implemented without early stopping and with the same number of (5) epochs as MDPO.
@@ F.1 Cartpole
Figure 8: Robust constrained MDP training development plots in the Cartpole domain, where each sample in the plot is based on 20 evaluations of the deterministic policy as it interacts with the adversarial environment from that iteration. The line and shaded area represent the mean and standard error across 10 seeds.
@@ F.2 Inventory Management
Figure 9: Constrained MDP training development plots in the Inventory Management domain, where each sample in the plot is based on 20 evaluations of the deterministic policy. The line and shaded area represent the mean and standard error across 10 seeds.
Figure 10: Robust constrained MDP training development plots in the Inventory Management domain, where each sample in the plot is based on 20 evaluations of the deterministic policy as it interacts with the adversarial environment from that iteration. The line and shaded area represent the mean and standard error across 10 seeds.
@@
Figure 11: Constrained MDP training development plots in the 3-D Inventory Management domain, where each sample in the plot is based on 20 evaluations of the deterministic policy. The line and shaded area represent the mean and standard error across 10 seeds.
Figure 12: Robust constrained MDP training development plots in the 3-D Inventory Management domain, where each sample in the plot is based on 20 evaluations of the deterministic policy as it interacts with the adversarial environment from that iteration. The line and shaded area represent the mean and standard error across 10 seeds.
@@ Appendix H Comparison of schedules
Tomar et al., (2022) propose either a fixed $\alpha$ or a linear increase across iterations to give a larger penalty at the end of the iterations (inspired by theoretical works, e.g. Beck and Teboulle, (2003)). A potential problem with such a linear scheme is that if the LTMA update is strong and the policy cannot take sufficiently large learning steps, it may sometimes lead to a gradual loss of performance and premature convergence, especially in the context of contrained optimisation where the objective is often subject to large changes due to constraints becoming active or inactive. As an alternative, we also consider a schedule where the penalty parameter schedule is restarted every time the LTMA update to $\xi$ is larger than a particular threshold. The schedule then follows the process $\alpha_{k}=\frac{1}{1-(k-k^{\prime})/(K-k^{\prime})}$, where $k^{\prime}$ indicates the restart time and $K$ is the final time. In the experiments, the restart is done every time any dimension $i$ of the update is larger than half the maximal distance from nominal, i.e. larger than $0.5\max_{\xi_{i}\in\mathcal{U}_{\xi_{i}}}|\xi_{i}-\bar{\xi}_{i}|$. A last included schedule is the geometrically decreasing schedule $\alpha_{k}=\alpha_{k-1}*\gamma$ based on the discount factor $\gamma$ as proposed in earlier work (Xiao,, 2022). We experiment with both epoch-based and batch-based schedules. While in the batch-based schedules we additionally experimented with the more aggressive update rule $\alpha_{k}=\alpha_{k-1}\frac{1-\gamma}{M}$ proposed in Wang et al., (2024). However, the results under the scheme were poor, and so these results are not included.

### v4

@@ Appendix E Algorithm implementation details
All the algorithms are implemented in Pytorch. The code for RMCPMD is based on the original implementation, which can be found at https://github.com/JerrisonWang/JMLR-DRPMD. The code for PPO is taken from the StableBaselines3 class. Both are modified to fit our purposes by allowing to turn off and on robust training and constraint-satisfaction. Similar to PPO, MDPO is also implemented following StableBaselines3 class conventions. All the environments are implemented in Gymnasium. For simplicity, the updates with LTMA are based on the original TMA code in https://github.com/JerrisonWang/JMLR-DRPMD, which performs pure Monte Carlo based TMA rather than additional function approximation as proposed in our Approximate TMA algorithm. For MDPO and PPO experiments, we use four parallel environments with four CPUs. Our longest experiments typically take no more than two hours to complete. The remainder of the section describes domain-specific hyperparameter settings.
@@ E.1 Hyperparameters for Cartpole experiments
Hyperparameters settings for the Carpole experiments can be found in 8. Settings for robust optimisation, including the policy architecture, transition kernel architecture, and TMA learning rate are taken from Wang et al. (2023) with the exception that we formulate a more challenging uncertainty set with a 5 times larger range. The policy learning rate and GAE lambda is typical for PPO based methods so we apply these for PPO and MDPO methods. The number of policy epochs and early stopping KL target is based for PPO methods on the standard repository for PPO-Lag (https://github.com/openai/safety-starter-agents/) and for MDPO it is based on the original paper’s settings (Tomar et al., 2022). PPO obtained better results without entropy regularisation and value function clipping on initial experiments, so for simplicity we disabled these for all algorithms. The batch size, multiplier initialisation, and multiplier learning rate and scaling are obtained via a limited tuning procedure. For the tuning, the batch size tried was in $\{100,200,400,1000,2000\}$ and the multiplier learning rate was set to $\{1\mathrm{e}^{-1},1\mathrm{e}^{-2},1\mathrm{e}^{-3},1\mathrm{e}^{-4}\}$ on partial runs. Then the use of a softplus transformation of the multiplier was compared to the direct (linear) multiplier, and for the linear multiplier we tested initialisations in $\{1,5,10\}$.
@@ E.2 Hyperparameters for Inventory Management experiments
Hyperparameters settings for the unidimensional Inventory Management (see Table 9) are similar to the Cartpole experiments. The key differences are the output layer, the discount factor, the transition kernel, and the TMA parameters, which follow settings of the domain in Wang et al. (2024). Based on initial tuning experiments with PPO and MDPO, the learning rates are set to $\eta=1\mathrm{e}^{-3}$ and $\eta_{\lambda}=1\mathrm{e}^{-2}$, a lower GAE lambda of 0.50 was chosen and PPO is implemented without early stopping and with the same number of (5) epochs as MDPO.
@@ F.1 Cartpole
Figure 8: RCMDP training development plots in the Cartpole domain, where each sample in the plot is based on 20 evaluations of the deterministic policy as it interacts with the adversarial environment from that iteration. The line and shaded area represent the mean and standard error across 10 seeds.
@@ F.2 Inventory Management
Figure 9: CMDP training development plots in the Inventory Management domain, where each sample in the plot is based on 20 evaluations of the deterministic policy. The line and shaded area represent the mean and standard error across 10 seeds.
Figure 10: RCMDP training development plots in the Inventory Management domain, where each sample in the plot is based on 20 evaluations of the deterministic policy as it interacts with the adversarial environment from that iteration. The line and shaded area represent the mean and standard error across 10 seeds.
@@
Figure 11: MDP training development plots in the 3-D Inventory Management domain, where each sample in the plot is based on 20 evaluations of the deterministic policy. The line and shaded area represent the mean and standard error across 10 seeds.
Figure 12: RCMDP training development plots in the 3-D Inventory Management domain, where each sample in the plot is based on 20 evaluations of the deterministic policy as it interacts with the adversarial environment from that iteration. The line and shaded area represent the mean and standard error across 10 seeds.
@@ Appendix H Comparison of schedules
Tomar et al. (2022) propose either a fixed $\alpha$ or a linear increase across iterations to give a larger penalty at the end of the iterations (inspired by theoretical works, e.g. Beck & Teboulle (2003)). A potential problem with such a linear scheme is that if the LTMA update is strong and the policy cannot take sufficiently large learning steps, it may sometimes lead to a gradual loss of performance and premature convergence, especially in the context of contrained optimisation where the objective is often subject to large changes due to constraints becoming active or inactive. As an alternative, we also consider a schedule where the penalty parameter schedule is restarted every time the LTMA update to $\xi$ is larger than a particular threshold. The schedule then follows the process $\alpha_{k}=\frac{1}{1-(k-k^{\prime})/(K-k^{\prime})}$, where $k^{\prime}$ indicates the restart time and $K$ is the final time. In the experiments, the restart is done every time any dimension $i$ of the update is larger than half the maximal distance from nominal, i.e. larger than $0.5\max_{\xi_{i}\in\mathcal{U}_{\xi_{i}}}|\xi_{i}-\bar{\xi}_{i}|$. A last included schedule is the geometrically decreasing schedule $\alpha_{k}=\alpha_{k-1}*\gamma$ based on the discount factor $\gamma$ as proposed in earlier work (Xiao, 2022). We experiment with both epoch-based and batch-based schedules. While in the batch-based schedules we additionally experimented with the more aggressive update rule $\alpha_{k}=\alpha_{k-1}\frac{1-\gamma}{M}$ proposed in Wang et al. (2024). However, the results under the scheme were poor, and so these results are not included.